the line parallel to 2x + 6y = -3 and passes through the y-axis at -3/4
Answer by mananth(16949)

2x + 6y = -3 and passes through the y-axis at -3/4
6y= -2x-3
y= -2/6 x -1/2
slope = m = -1/3
the parallel line will have same slope
passing through (0,-3/4)
So the equation is y = (-1/3)x-(3/4)